OpenStudy (anonymous):

Solve for x: 3x − 5 = 2x + 6

OpenStudy (anonymous):

Subtract 2x from each side

OpenStudy (anonymous):

Then add 5 to each side.

OpenStudy (anonymous):

3x-5=2x+6 3x=2x+11 x=11

OpenStudy (anonymous):

@hheston, please dont give out the answer, but help the OP in answering the question by themselves.